Find the probability of the following.


A. P(z < 2.45)


B. P(1.2<z<1.4)

1
Expert's answer
2022-02-21T15:53:56-0500

"P(Z<2.45)=0.9929."


"P(1.2<Z<1.4)=P(Z<1.4)-P(Z<1.2)=0.9192-0.8849=0.0343."